I need help with a few problems. Could you walk me through step by step on how to do them?
1) 5x^2 -72=0
Sorry to bother you with this but i really need help.
Thank you,
Jed LaCoste

Answer by richwmiller(17219)

some website rules
one problem at a time
no similar problems
limit 4 problems daily
and everyone here really needs help. What makes you more important than the next?
5x^2=72
x^2=72/5
x=+\-sqrt(72/5)
x=+\-6*sqrt(2/5)